# Demystifying Deep Search: A Holistic Evaluation with Hint-free Multi-Hop Questions and Factorised Metrics

### v2

@@ 1 Introduction
Web agents—autonomous systems that navigate and extract information from the internet—are becoming essential for extending language models beyond their parametric knowledge, as they tackle complex information-seeking tasks by integrating external search with internal knowledge, aggregating evidence from multiple sources, and synthesising it into coherent answers  Nakano et al. (2021); Ferrag et al. (2025). Among the evaluation scenarios for these systems Deep search is a central and foundational task: it requires uncovering hidden facts or entities through multi-step reasoning, noise filtering, and tackling the persistent “I can’t find it” challenge that even expert human searchers face. Recent efforts have sought to advance performance by designing stronger agent systems Li et al. (2025b; c); Sun et al. (2025); Jiang et al. (2025); Li et al. (2025a), while others have introduced new benchmarks to better characterise its nature Du et al. (2025); Wong et al. (2025).
@@
However, we identify a critical yet overlooked dimension in current deep search evaluation: the presence of various forms of hinting in question formulations, which fundamentally alters the nature of the task. As shown in Figure 1, classical multi-hop QA datasets such as HotpotQA Yang et al. (2018a) exhibit Path-Hinting (PH), where questions explicitly narrate the reasoning chain—for example, “Who is the husband of the stepmother of the brother of Kane Cornes?” directly instructs the agent to sequentially find the brother, then the stepmother, then her husband, reducing reasoning to execution. More recent benchmarks, including BrowseComp Wei et al. (2025) and WebShaper Tao et al. (2025), attempt to mitigate this issue through Specification-Hinting (SH): rather than naming the target entity, questions describe it indirectly through multiple attributes. For instance, instead of asking directly about “Graham Cornes,” the query specifies “Which radio presenter at 5AA, former footballer, at least 20 years senior to his wife who was a 2007 Labor candidate?”—requiring constraint satisfaction through filtering rather than genuine exploratory reasoning.
Such guided reasoning bypasses models’ need to autonomously search for information and construct reasoning paths, failing to test a critical capability of web-agents: the capacity to autonomously discover which connections matter, generate hypotheses about possible reasoning trajectories, and adaptively explore the information space without guidance. When agents are provided either an explicit path (PH) or a unique signature (SH), they operate with strong scaffolding that is rarely available in real-world scenarios. Furthermore, existing evaluations suffer from a critical limitation: they typically report only aggregate pass rates, collapsing diverse failure modes into a single metric. This obscures crucial distinctions—an agent that searches exhaustively but fails to connect evidence exhibits fundamentally different limitations than one that gives up prematurely or misuses its parametric knowledge. Without understanding these failure modes, it becomes hard to diagnose system weaknesses or guide improvements.
@@
In this work, we introduce WebDetective, a benchmark that rethinks hint-free deep search evaluation through the co-design of questions and their environment. We construct Hint-Free (HF) Multi-Hop questions that avoid both path narration and attribute fingerprints—for example, “Who is the father of Kane Cornes?”—requiring agents to autonomously discover relevant contexts and reasoning chains. A controlled Wikipedia sandbox prevents shortcuts by revealing information only when the correct reasoning path is followed, ensuring that intermediate links must be discovered sequentially. This co-design guarantees that successful answers reflect genuine multi-hop reasoning and enables a two-level evaluation framework with precise attribution of failure modes: we can separate knowledge sufficiency (from search or parametric sources) from generation quality, and determine whether an agent failed by halting too early, missing connections, or refusing appropriately when evidence was lacking. Such fine-grained diagnostics are possible only because the benchmark enforces a unique path to the correct answer and records exactly how far along that path each agent progresses.
Additionally, we further design an agentic workflow baseline EvidenceLoop that explicitly incorporates context retention, memory management, and verification steps, offering a first attempt at addressing the unique challenges posed by hint-free deep search. Through our diagnostic evaluation framework, we uncover fundamental brittleness in current systems when reasoning paths must be discovered rather than given, exposing critical gaps between existing capabilities and the requirements of genuine autonomous deep search.
@@ 2.1 Hint-Free Multi-Hop Question Answering
We define two prevalent hint types in existing multi-hop QA benchmarks Yang et al. (2018b); Chen et al. (2019); Wei et al. (2025):
@@ 2.2 The Co-Design Principle
To address this, we introduce a co-designed evaluation system that jointly constructs questions and environment to enforce reasoning-path discovery. Our controlled sandbox applies selective entity masking: for a chain $v_{0}\rightarrow v_{1}\rightarrow\cdots\rightarrow v_{n}$ (where $v_{0}$ appears in the question and $v_{n}$ yields the answer), each intermediate $v_{i}$ is masked everywhere except on $v_{i-1}$’s page:
@@
This enforces strict sequential access, eliminating shortcuts. For example, in “Who is the father of Kane Cornes?”, “Chad Cornes” appears only on Kane’s page, “Nicole Cornes” only on Chad’s, and “Graham Cornes” only on Nicole’s, requiring the chain Kane $\rightarrow$ Chad $\rightarrow$ Nicole $\rightarrow$ Graham. Success thus guarantees full path discovery, while failures can be precisely diagnosed—whether the agent failed to explore, missed a relevant step, or struggled with synthesis. This design turns multi-hop QA evaluation from probabilistic judgment into deterministic verification.
@@ 3 A Baseline Attempt: The EvidenceLoop Framework
To address the unique challenges posed by hint-free multi-hop reasoning, we develop EvidenceLoop, an agentic workflow baseline that explicitly incorporates context retention, memory management, and verification steps to maintain reasoning coherence across extended search trajectories. Unlike standard ReAct implementations that can lose track of evidence across many search iterations, our workflow introduces structured mechanisms for tracking discovered entities, maintaining evidence chains, and verifying reasoning paths before answer generation. We provide a detailed description of the EvidenceLoop architecture, including its controller configuration, memory modules, and verification procedures in Appendix D.
@@ 4 Experiments
We evaluate 25 state-of-the-art models with ReAct-style tool use, including those from OpenAI, Anthropic, Google, xAI, Alibaba, ByteDance, Zhipu AI, Moonshot AI, and High-Flyer. All models interleave reasoning, search, and observations in our controlled Wikipedia sandbox on WebDetective with 200 hint-free multi-hop questions (2–4 hops), under limits of 40 tool calls and a 32K-token context. Unless noted, decoding uses temperature=0.6, top_p=0.95. We report six metrics: (1) Knowledge Score, knowledge sufficiency; (2) Search Score, retrieval effectiveness; (3) Generation Score, weighted F1 of Good Refusal and Knowledge Utilization; (4) Good Refusal F1; (5) Knowledge Utilization F1; and (6) Pass@1. We additionally analyse Forget and Lead-astray to probe knowledge degradation (section 4.2.2). For our EvidenceLoop, we set breadth=3, iteration=3. Comprehensive results appear in table 1.
@@ 4.1 Main Results
Table 1: Comparison of 25 state-of-the-art models with ReAct-style tool use capabilities. Metrics cover Knowledge Discovery (Knowledge Sufficiency, Search Score), Generation Quality (Generation Score, Good Refusal F1, Knowledge Utilisation F1), Knowledge Degradation (Forget, Lead-astray), and Pass@1. Bold values denote best results: higher is better for Knowledge Discovery, Generation Quality, and Pass@1, while lower indicates greater robustness for Knowledge Degradation.
@@
GPT-OSS-120B OpenAI (2025b) 16.00 23.50 2.75 23.59 10.73 100.00 0.00 24.00
o3-Mini OpenAI (2025c) 48.50 57.00 9.10 21.05 16.48 46.39 42.27 21.50
o4-Mini OpenAI (2025d) 68.00 72.00 12.69 19.75 17.56 27.94 59.56 21.00
o3 OpenAI (2025c) 70.00 76.00 18.29 3.29 48.97 24.29 24.29 53.50
o3-Pro OpenAI (2025c) 71.00 78.00 20.86 9.37 49.40 21.83 25.35 56.00
GPT-5-Chat OpenAI (2025a) 58.00 59.50 15.74 26.23 28.05 47.41 31.90 29.50
GPT-5 OpenAI (2025a) 79.00 80.00 23.21 8.89 49.58 17.72 32.91 50.50
@@
Claude-Sonnet-4-Think Anthropic (2025) 66.50 73.50 26.19 34.59 44.19 45.11 21.80 38.50
Claude-Opus-4-Think Anthropic (2025) 68.00 73.50 21.00 30.53 31.23 43.38 32.35 29.00
Claude-Opus-4.1 Anthropic (2025) 74.00 76.50 28.53 28.57 48.54 27.03 31.08 44.50
@@
Gemini-2.5-Flash-Think Google DeepMind (2025) 59.00 64.50 16.79 40.56 16.35 57.63 35.59 17.50
Gemini-2.5-Pro Google DeepMind (2025) 65.50 73.00 11.64 10.87 24.68 44.27 35.11 28.50
@@
Qwen3-30B-Think Yang et al. (2025) 56.50 59.00 7.25 12.51 13.16 79.65 16.81 12.50
Qwen3-235B-Think Yang et al. (2025) 72.50 72.00 11.15 6.56 24.19 63.45 19.31 21.50
Tongyi-DeepResearch Tongyi DeepResearch Team (2025) 53.50 57.50 4.20 0.00 15.69 43.93 41.12 18.50
@@
Doubao-1.6-Flash ByteDance Seed Team (2025) 54.50 57.50 20.00 53.95 19.46 68.81 21.10 13.50
Doubao-1.6-Think ByteDance Seed Team (2025) 64.00 68.50 19.24 42.03 18.11 49.22 39.84 16.00
@@
GLM-4.5-Air-Inner Zhipu AI Team (2025) 55.50 60.50 12.31 26.39 17.97 44.14 40.54 19.00
GLM-4.5-Inner Zhipu AI Team (2025) 63.50 67.50 22.19 34.79 35.09 25.98 40.16 33.50
@@
Kimi-K2-0711 Moonshot AI (2025) 54.50 59.00 9.72 16.36 19.31 43.12 36.70 23.50
Kimi-K2-0905 Moonshot AI (2025) 53.00 55.00 13.17 28.79 20.89 49.06 33.96 24.00
@@
DeepSeek-R1 DeepSeek-AI et al. (2025) 61.50 65.50 10.57 18.81 15.55 37.40 51.22 20.00
DeepSeek-V3.1  DeepSeek-AI et al. (2024) 61.50 56.50 13.62 27.97 16.34 44.72 44.72 17.00
DeepSeek-V3.1-Terminus  DeepSeek-AI et al. (2024) 55.50 58.50 16.31 36.49 22.23 28.83 50.45 24.50

### v3

@@ 1 Introduction
Web agents—autonomous systems that navigate and extract information from the internet—are becoming essential for extending language models beyond their parametric knowledge, as they tackle complex information-seeking tasks by integrating external search with internal knowledge, aggregating evidence from multiple sources, and synthesising it into coherent answers  nakano2021webgpt; ferrag2025llmreasoningautonomousai. Among the evaluation scenarios for these systems Deep search is a central and foundational task: it requires uncovering hidden facts or entities through multi-step reasoning, noise filtering, and tackling the persistent “I can’t find it” challenge that even expert human searchers face. Recent efforts have sought to advance performance by designing stronger agent systems li2025searcho1agenticsearchenhancedlarge; li2025webthinkerempoweringlargereasoning; sun2025zerosearchincentivizesearchcapability; jiang2025s3dontneeddata; li2025websailornavigatingsuperhumanreasoning, while others have introduced new benchmarks to better characterise its nature du2025deepresearchbenchcomprehensivebenchmark; wong2025widesearchbenchmarkingagenticbroad.
However, we identify a critical yet largely overlooked issue in current deep-search evaluation: many benchmarks contain forms of hinting that guide models toward the answer and remove the need for genuine autonomous reasoning. Such hints bypass a core capability of web agents—the ability to discover which connections matter, generate hypotheses about plausible reasoning trajectories, and adaptively explore the information space without guidance.
@@
As illustrated in Figure 1, classical multi-hop QA datasets such as HotpotQA yang2018hotpotqadatasetdiverseexplainable exhibit Path-Hinting (PH), where the question explicitly encodes the reasoning chain (e.g., “Who is the husband of the stepmother of the brother of Kane Cornes?”), reducing the task to executing a prescribed sequence. More recent benchmarks, including BrowseComp wei2025browsecomp and WebShaper tao2025webshaperagenticallydatasynthesizing, introduce Specification-Hinting (SH), describing the target entity through multiple attributes rather than naming it directly (e.g., “Which radio presenter at 5AA, former footballer, at least 20 years senior to his wife…?”). Although less explicit, this still provides a unique signature that primarily requires constraint filtering rather than exploratory reasoning. In both cases, agents operate with scaffolding rarely available in real-world scenarios, limiting the evaluation of true autonomous search.
Furthermore, existing evaluations suffer from another critical limitation: they typically report only an aggregate pass rate, collapsing diverse failure modes into a single number. This masks essential differences—an agent that searches thoroughly but fails to synthesise evidence has fundamentally different weaknesses from one that halts early, misuses its parametric knowledge, or refuses to reason when it should proceed. Likewise, when evidence is insufficient, the model should recognise and signal its knowledge gap rather than produce unsupported answers. Without disentangling these behaviours, it is difficult to diagnose system deficiencies or to meaningfully guide future improvements.
@@
In this work, we introduce WebDetective, a benchmark that rethinks deep-search evaluation through the co-design of hint-free questions and a controlled environment. We construct Hint-Free (HF) Multi-Hop questions that avoid both path narration and attribute fingerprints while retaining a unique answer (e.g., “Who is the father of Kane Cornes?”). These questions are paired with a controlled Wikipedia sandbox containing a reference reasoning path, enabling fine-grained diagnostics. To prevent shortcutting, each node in the reference chain is masked so that the next entity can be accessed only through its predecessor. This design preserves the ability to compare any valid reasoning strategy against a canonical shortest search-only path—reflecting real settings with multiple possible routes—while decoupling evaluation into search ability, information synthesis, and final task success. Additionally, we further design an agentic workflow baseline EvidenceLoop that explicitly incorporates context retention, memory management, and verification steps, offering a first attempt at addressing the unique challenges posed by hint-free deep search. Through our diagnostic evaluation framework, we uncover fundamental brittleness in current systems when reasoning paths must be discovered rather than given, exposing critical gaps between existing capabilities and the requirements of genuine autonomous deep search.
Based on the diagnostic framework, we evaluate 25 representative state-of-the-art models and observe a consistent pattern: when deprived of hints, even the strongest systems succeed only on a limited questions, and their failures differ markedly from what existing benchmarks suggest. Many models retrieve ample evidence but fail to synthesise it into a coherent reasoning chain; others prematurely abandon search or generate unsupported answers despite clear knowledge gaps, revealing weak calibrated refusal. These behaviours stand in sharp contrast to their strong performance on conventional multi-hop datasets and highlight the need to evaluate the ability to discover reasoning paths rather than merely execute them.
Our contributions are as follows:
  • We introduce WebDetective, the first benchmark to systematically eliminate both path hinting and specification hinting while preserving a controlled environment that supports fine-grained diagnostic analysis.
  • We propose a comprehensive evaluation framework that disentangles search sufficiency, information synthesis, and refusal behaviour, enabling precise identification of failure modes in deep-search agents.
  • We develop EvidenceLoop, an agentic workflow with explicit context retention, memory management, and verification mechanisms, showing that our benchmark not only reveals system weaknesses but also informs concrete architectural improvements.
@@ 2.1 Hint-Free Multi-Hop Question Answering
We define two prevalent hint types in existing multi-hop QA benchmarks yang2018hotpotqa; chen2019multihop; wei2025browsecomp:
@@ 2.2 The Co-Design Principle
This masking eliminates corpus-based shortcuts—agents cannot find the answer through spurious co-occurrences or unrelated contexts—while still accommodating diverse valid reasoning strategies. The ground-truth chain serves as a reference path (the shortest sufficient route using the corpus alone), but our evaluation framework does not enforce strict adherence to this path. To comprehensively assess whether an agent’s search is sufficient and efficient, we employ a multi-faceted evaluation:
(1) Shortest Reference Chain Coverage. We first check whether all hops in the reference chain were visited. For any missing evidence, we probe the model’s parametric knowledge with targeted queries (e.g., “Kane Cornes has brother    ?”). If visited evidence plus verified parametric knowledge covers the complete chain, the instance achieves knowledge sufficiency.
(2) Alternative Path Recognition. Agents may discover valid reasoning routes different from the reference chain. We collect all evidence the agent actually visited during search and feed this clean context to an LLM judge. If the judge can correctly answer the question from this visited evidence alone, the instance also achieves knowledge sufficiency—regardless of whether the reference chain was followed. Consider one real example from our benchmark: “On which compilation album does the song ‘Loneliness of a Middle Distance Runner’ appear?” The reference path follows: Novel (The Loneliness of the Long-Distance Runner) $\rightarrow$ Single (Jonathan David) $\rightarrow$ Compilation (Push Barman to Open Old Wounds), tracing through the single that featured the song as a B-side. However, an agent might instead traverse Novel $\rightarrow$ Belle and Sebastian (the band mentioned as adapting the title) $\rightarrow$ the band’s discography page listing “Push Barman to Open Old Wounds” as a B-sides compilation—reaching the same answer through the artist’s discography rather than the specific single.
(3) Efficient Parametric Knowledge Use. Our Search Score further credits agents that efficiently combine search with parametric knowledge. When an agent correctly answers the question using fewer (or equal) hops than the reference path while actively performing search, this demonstrates intelligent leveraging of internal knowledge to shortcut the reasoning process, and receives additional credit.
Returning to our running example, “Who is the father of Kane Cornes?”, the reference path is Kane $\rightarrow$ Chad (brother) $\rightarrow$ Nicole (stepmother) $\rightarrow$ Graham (husband). An agent achieves knowledge sufficiency by: (a) visiting all four pages, (b) visiting partial pages while parametric probes confirm missing knowledge, or (c) discovering alternative evidence that supports the correct answer. An agent that correctly answers using only two hops (e.g., Kane $\rightarrow$ Chad, then leveraging parametric knowledge about Chad’s family) receives Search Score credit for efficient reasoning. This design eliminates unearned shortcuts while rewarding genuine multi-hop reasoning through any valid strategy.
Lastly, while we instantiate this benchmark using Wikipedia as a controlled testbed, the co-design methodology is not inherently tied to this corpus. The framework generalises to any domain satisfying three conditions: (1) a text corpus with factual content, (2) an entity-level graph or link structure over that corpus, and (3) the ability to mask entity mentions along chosen paths. Given an appropriate dataset meeting these requirements, the same evaluation pipeline—path construction, selective masking, and factorised metrics—can be directly applied to news archives, scientific repositories, enterprise knowledge bases, or curated web snapshots, extending the diagnostic value beyond the current Wikipedia instantiation.
@@ 2.4 Dataset Construction
Domain Coverage. The final dataset spans diverse domains to ensure broad evaluation coverage: Film, Television & Theatre (20.0%), History, Politics & Military (18.0%), Geography & Landmarks (16.0%), Music (13.5%), Sports (11.5%), Biography & Education (8.5%), Literature & Pop Culture (6.5%), and Science & Technology (6.0%). This distribution ensures WebDetective is not skewed toward any single relation type but covers reasoning patterns essential for evaluating multi-hop capabilities.
@@ 3 A Baseline Attempt: The EvidenceLoop Framework
To address the unique challenges posed by hint-free multi-hop reasoning, we develop EvidenceLoop, an agentic workflow baseline that explicitly incorporates context retention, memory management, and verification steps to maintain reasoning coherence across extended search trajectories. Unlike standard ReAct implementations that can lose track of evidence across many search iterations, our workflow introduces structured mechanisms for tracking discovered entities, maintaining evidence chains, and verifying reasoning paths before answer generation.
@@
Positioning Relative to Existing Frameworks. EvidenceLoop is a diagnostically-derived workflow designed to address the specific failure modes revealed by WebDetective. Table 1 summarises the key architectural differences from standard approaches. We provide a more detailed description of the EvidenceLoop architecture, including its controller configuration, memory modules, and verification procedures in Appendix E.
Table 1: Architectural comparison of EvidenceLoop with existing agentic frameworks (Standard ReAct and ReAct with Reflection).
Aspect Standard ReAct ReAct + Self-Reflection EvidenceLoop (Ours)
Controller Loop Single thought-tool-observation Same + occasional reflection Two-phase: exploration then verification
Memory Flat dialogue history Same + reflection messages Structured buffer: (entity, snippet, source) tuples
Verification Often none Critiques not tied to evidence Explicit loop using only evidence buffer
Breadth & Iterations Single trajectory Single trajectory Parallel trajectories + explore-aggregate cycles
Refusal Behaviour Implicit, rarely triggered Loosely specified Explicitly tied to verification: incomplete $\rightarrow$ refuse
@@ 4 Experiments
We evaluate 25 state-of-the-art models with ReAct-style tool use, including those from OpenAI, Anthropic, Google, xAI, Alibaba, ByteDance, Zhipu AI, Moonshot AI, and High-Flyer. All models interleave reasoning, search, and observations in our controlled Wikipedia sandbox on WebDetective with 200 hint-free multi-hop questions (2–4 hops), under limits of 40 tool calls and a 32K-token context. Unless noted, decoding uses temperature=0.6, top_p=0.95. We report six metrics: (1) Knowledge Score, knowledge sufficiency; (2) Search Score, retrieval effectiveness; (3) Generation Score, weighted F1 of Good Refusal and Knowledge Utilization; (4) Good Refusal F1; (5) Knowledge Utilization F1; and (6) Pass@1. We additionally analyse Forget and Lead-astray to probe knowledge degradation (Section 4.2.2). For our EvidenceLoop, we use DeepSeek-R1 as the base model and set breadth=3, iteration=3. Comprehensive results appear in Table 2.
@@ 4.1 Main Results
Table 2: Comparison of 25 state-of-the-art models with ReAct-style tool use capabilities. Metrics cover Knowledge Discovery (Knowledge Sufficiency, Search Score), Generation Quality (Generation Score, Good Refusal F1, Knowledge Utilisation F1), Knowledge Degradation (Forget, Lead-astray), and Pass@1. Bold values denote best results: higher is better for Knowledge Discovery, Generation Quality, and Pass@1, while lower indicates greater robustness for Knowledge Degradation. EvidenceLoop uses DeepSeek-R1 as its base model; shaded rows highlight this comparison pair.
@@
GPT-OSS-120B openai2025gptoss 16.00 23.50 2.75 23.59 10.73 100.00 0.00 24.00
o3-Mini openai2025o3 48.50 57.00 9.10 21.05 16.48 46.39 42.27 21.50
o4-Mini openai2025o4 68.00 72.00 12.69 19.75 17.56 27.94 59.56 21.00
o3 openai2025o3 70.00 76.00 18.29 3.29 48.97 24.29 24.29 53.50
o3-Pro openai2025o3 71.00 78.00 20.86 9.37 49.40 21.83 25.35 56.00
GPT-5-Chat openai2025gpt5 58.00 59.50 15.74 26.23 28.05 47.41 31.90 29.50
GPT-5 openai2025gpt5 79.00 80.00 23.21 8.89 49.58 17.72 32.91 50.50
@@
Claude-Sonnet-4-Think anthropic2025claude4 66.50 73.50 26.19 34.59 44.19 45.11 21.80 38.50
Claude-Opus-4-Think anthropic2025claude4 68.00 73.50 21.00 30.53 31.23 43.38 32.35 29.00
Claude-Opus-4.1 anthropic2025claude4 74.00 76.50 28.53 28.57 48.54 27.03 31.08 44.50
@@
Gemini-2.5-Flash-Think google2025gemini 59.00 64.50 16.79 40.56 16.35 57.63 35.59 17.50
Gemini-2.5-Pro google2025gemini 65.50 73.00 11.64 10.87 24.68 44.27 35.11 28.50
@@
Qwen3-30B-Think qwen2025qwen3 56.50 59.00 7.25 12.51 13.16 79.65 16.81 12.50
Qwen3-235B-Think qwen2025qwen3 72.50 72.00 11.15 6.56 24.19 63.45 19.31 21.50
Tongyi-DeepResearch alibaba2025tongyi 53.50 57.50 4.20 0.00 15.69 43.93 41.12 18.50
@@
Doubao-1.6-Flash bytedance2025doubao 54.50 57.50 20.00 53.95 19.46 68.81 21.10 13.50
Doubao-1.6-Think bytedance2025doubao 64.00 68.50 19.24 42.03 18.11 49.22 39.84 16.00
@@
GLM-4.5-Air-Inner zhipuai2025glm 55.50 60.50 12.31 26.39 17.97 44.14 40.54 19.00
GLM-4.5-Inner zhipuai2025glm 63.50 67.50 22.19 34.79 35.09 25.98 40.16 33.50
@@
Kimi-K2-0711 moonshot2025kimi 54.50 59.00 9.72 16.36 19.31 43.12 36.70 23.50
Kimi-K2-0905 moonshot2025kimi 53.00 55.00 13.17 28.79 20.89 49.06 33.96 24.00
@@
DeepSeek-R1 deepseek2025r1 61.50 65.50 10.57 18.81 15.55 37.40 51.22 20.00
DeepSeek-V3.1  deepseek2025v3 61.50 56.50 13.62 27.97 16.34 44.72 44.72 17.00
DeepSeek-V3.1-Terminus  deepseek2025v3 55.50 58.50 16.31 36.49 22.23 28.83 50.45 24.50
@@
EvidenceLoop demonstrates targeted improvements.
Comparing EvidenceLoop to its base model DeepSeek-R1 reveals meaningful gains from our diagnostically-derived design: Pass@1 improves from 20.0% to 25.0% (+25% relative), Generation Score from 10.57% to 12.61% (+19% relative), and most notably, Knowledge Utilization F1 from 15.55% to 23.79% (+53% relative). The substantial improvement in Knowledge Utilization demonstrates that our structured evidence buffer and verification loop directly address the synthesis bottleneck that WebDetective identifies as a critical failure mode. While absolute gains are moderate, these improvements validate that the failure modes exposed by our diagnostic framework can guide concrete architectural improvements. We conduct additional experiments across different base models on EvidenceLoop and investigate the effectiveness of each component in Appendix E.